# FedSRD: Sparsify-Reconstruct-Decompose for Communication-Efficient Federated Large Language Models Fine-Tuning

### v1

@@ Abstract.
The current paradigm of training large language models (LLMs) on publicly available Web data is becoming unsustainable, with high-quality data sources in specialized domains nearing exhaustion. Federated Learning (FL) emerges as a practical solution for the next generation of AI on a decentralized Web, enabling privacy-preserving collaborative fine-tuning by leveraging private data distributed across a global client base. While Low-Rank Adaptation (LoRA) is the standard for efficient fine-tuning, its application in federated settings presents a critical challenge: communication overhead remains a significant bottleneck across the Web’s heterogeneous network conditions. The structural redundancy within LoRA parameters not only incurs a heavy communication burden but also introduces conflicts when aggregating client updates. To address this, we propose FedSRD, a Sparsify-Reconstruct-Decompose framework designed for communication-efficient FL. We first introduce an importance-aware sparsification method that preserves the structural integrity of LoRA updates to reduce the uploaded parameter count. The server then reconstructs and aggregates these updates in a full-rank space to mitigate conflicts. Finally, it decomposes the global update into a sparse low-rank format for broadcast, ensuring a symmetrically efficient cycle. We also propose an efficient variant, FedSRD-e, to reduce computational overhead. Experimental results on 10 benchmarks demonstrate that our framework significantly reduces communication costs by up to 90% while even improving model performance on heterogeneous client data.
@@ 1. Introduction
Large Language Models (LLMs), pre-trained on vast web-scale corpora, have demonstrated remarkable generalist capabilities (Ouyang et al., 2022; Kaplan et al., 2020). However, realizing their full potential for specialized downstream applications requires fine-tuning on diverse, high-quality, domain-specific data. Such data is often proprietary and distributed across numerous private sources, resulting in fragmented ”data silos.” As centralizing this sensitive data is frequently impractical due to privacy regulations and cost constraints, Federated Learning (FL) offers a compelling solution, enabling multiple clients to collaboratively fine-tune a global shared LLM without direct data exchange. This paradigm is central to building the next generation of intelligent, decentralized Web applications.
@@
Despite the efficiency of LoRA, the iterative communication of its parameters remains a substantial overhead. This becomes a critical bottleneck when deployed across the highly heterogeneous and unpredictable network conditions characteristic of the global Web and mobile systems (Dorfman et al., 2023; Mundt, 2004). High communication costs can exclude clients with limited bandwidth, such as those in rural areas or on mobile devices, undermining data diversity and posing fairness and inclusivity challenges for the global Web ecosystem. While sparsification is a common strategy for reducing communication, conventional methods are ill-adapted for federated LoRA fine-tuning on non-IID data. As shown in Figure 2, naïve approaches like magnitude-based pruning (Aji and Heafield, 2017; Yadav et al., 2023; Deep et al., 2024) disrupt the critical structural relationship between the LoRA matrices $A$ and $B$, while random sparsification (Yu et al., 2024) also leads to performance degradation at high sparsity ratios. Crucially, both methods fail to preserve client-specific knowledge and exacerbate conflicts during server-side aggregation under data heterogeneity. This results in an undesirable trade-off: reduced communication at the expense of significant performance loss. This leads us to our central research question:
@@
To address this question, we introduce FedSRD, a novel Sparsify-Reconstruct-Decompose framework for communication-efficient federated LLM fine-tuning. Our framework is designed to maintain high model performance through three sequential operations.
@@ 2.3. Communication-Efficient Federated LLMs Finetuning
Communication overhead remains a primary bottleneck in FL systems (Yan et al., 2025, 2024; Sattler et al., 2019), a problem that is exacerbated when fine-tuning LLMs over the public internet. While PEFT methods provide a baseline level of efficiency, further optimizations are crucial. Structured methods impose constraints on the LoRA updates. For instance, FFA-LoRA (Sun et al., 2024) halves the communication cost by transmitting only the $B$ matrices. Building on this, FedSVD (Lee et al., 2025) iteratively re-initializes LoRA $A$ matrices. Fed-SB (Singhal et al., 2025) freezes the LoRA matrices entirely and instead inserts a small, trainable square matrix for aggregation. While these approaches reduce communication, they often do so at the cost of model performance by constraining the expressive capacity of LoRA. Other lines of work focus on generic compression. Quantization methods (Aji and Heafield, 2017; Konečnỳ et al., 2016) represent parameters with fewer bits, but they often fail to achieve high compression ratios and can lead to significant accuracy degradation, particularly in non-IID settings. Sparsification methods, which transmit only a subset of non-zero parameters, can achieve much higher compression ratios (Aji and Heafield, 2017; Sahu et al., 2021). Yet, traditional approaches are not directly applicable to LoRA. Magnitude-based pruning (Aji and Heafield, 2017) is suboptimal because it ignores the coupled relationship between the two LoRA matrices. Similarly, random sparsification (Yang et al., 2025; Yu et al., 2024) fails to preserve critical parameters, resulting in inferior performance at high sparsity ratios.
@@ 4.3. Alternative Decomposition
After aggregating in the full-rank space, the server must decompose the result back into low-rank matrices for the clients. Previous works (Bai et al., 2024; Lee et al., 2025) often use Singular Value Decomposition (SVD) to construct new dense matrices $B^{t}$ and $A^{t}$. However, transmitting these dense matrices would create a severe download bottleneck, negating the benefit of a sparse upload. Naively sparsifying the full state matrices $B^{t}$ and $A^{t}$ would damage the learned model, and one cannot simply compute an update as $B^{t}-B^{t-1}$ because it does not correctly approximate the global update ($B^{t}A^{t}-B^{t-1}A^{t-1}\neq(B^{t}-B^{t-1})A^{t-1}$). This motivates a novel decomposition method that can produce a sparse update for download.
@@ 5.3. Computation Time Analysis
We measured the wall-clock time required for a single FL round in our simulation environment for both the Llama3.2-3B and Qwen2-7B models, with results reported in Table 7. Among the baselines, FedAvg provides a baseline for time consumption. FedProx requires slightly more time due to its proximal term, whereas FFA-LoRA is faster as it only trains and aggregates half of the LoRA matrices. FlexLoRA consumes more time performing SVD for matrix decomposition, and FedICU is the most time-intensive due to its computationally inefficient disentangled aggregation. For our methods, while the full FedSRD requires more time for its $r$-rank SVD projection, our efficient variant, FedSRD-e, introduces negligible additional computation compared to FedAvg, while exhibiting performance second only to the full FedSRD. This result highlights that our FedSRD-e variant achieves a Pareto-optimal balance across model performance, communication efficiency, and time consumption, making it a highly practical and scalable solution for real-world federated LLM fine-tuning.
@@ 6. Conclusion
In this paper, we address the critical challenge of communication overhead in federated LoRA fine-tuning for LLMs. We propose FedSRD, a novel Sparsify-Reconstruct-Decompose framework for communication-efficient FL. Its core components include a structure-aware, importance-based sparsification strategy on the client side; a conflict-mitigating reconstruction and aggregation step in the full-rank space on the server side; and a novel alternating decomposition method to produce sparse global updates for efficient download. We also introduce FedSRD-e, a computationally light variant for resource-constrained environments. Through extensive experiments on diverse benchmarks, we demonstrated that FedSRD reduces communication costs by over 90% while simultaneously improving model performance compared to state-of-the-art methods. Our work presents a practical and high-performing solution that significantly lowers the barrier for scalable, collaborative AI development on the decentralized Web.

### v2

@@ Abstract.
The current paradigm of training large language models (LLMs) on publicly available Web data is becoming unsustainable, with high-quality data sources in specialized domains nearing exhaustion. Federated Learning (FL) emerges as a practical solution for the next generation of AI on a decentralized Web, enabling privacy-preserving collaborative fine-tuning by leveraging private data distributed across a global client base. While Low-Rank Adaptation (LoRA) is the standard for efficient fine-tuning, its application in federated settings presents a critical challenge: communication overhead remains a significant bottleneck across the Web’s heterogeneous network conditions. The structural redundancy within LoRA parameters not only incurs a heavy communication burden but also introduces conflicts when aggregating client updates. To address this, we propose FedSRD, a Sparsify-Reconstruct-Decompose framework designed for communication-efficient federated LLMs fine-tuning. We first introduce an importance-aware sparsification method that preserves the structural integrity of LoRA updates to reduce the uploaded parameter count. The server then reconstructs and aggregates these updates in a full-rank space to mitigate conflicts. Finally, it decomposes the global update into a sparse low-rank format for broadcast, ensuring a symmetrically efficient cycle. We also propose an efficient variant, FedSRD-e, to reduce computational overhead. Experimental results on 10 benchmarks demonstrate that our framework significantly reduces communication costs by up to 90% while even improving model performance on heterogeneous client data.
@@ 1. Introduction
Large Language Models (LLMs), pre-trained on vast web-scale corpora, have demonstrated remarkable generalist capabilities (Ouyang et al., 2022; Kaplan et al., 2020). However, realizing their full potential for specialized downstream applications requires fine-tuning on diverse, high-quality, domain-specific data. Such data is often proprietary and distributed across numerous private sources, resulting in fragmented ”data silos”. As centralizing this sensitive data is frequently impractical due to privacy regulations and cost constraints, Federated Learning (FL) offers a compelling solution, enabling multiple clients to collaboratively fine-tune a global shared LLM without direct data exchange. This paradigm is central to building the next generation of intelligent, decentralized Web applications (Sani et al., [n. d.]; Wu et al., 2025).
@@
Despite the efficiency of LoRA, the iterative communication of its parameters remains a substantial overhead. This becomes a critical bottleneck when deployed across the highly heterogeneous and unpredictable network conditions characteristic of the global Web and mobile systems (Dorfman et al., 2023; Mundt, 2004). High communication costs can exclude clients with limited bandwidth, such as those in rural areas or on mobile devices, undermining data diversity and posing fairness and inclusivity challenges for the global Web ecosystem. While sparsification is a common strategy for reducing communication, conventional methods are ill-adapted for federated LoRA fine-tuning on non-IID data. As shown in Figure 2, naïve approaches like magnitude-based pruning (Aji and Heafield, 2017) disrupt the critical structural relationship between the LoRA matrices $A$ and $B$, while random sparsification (Yu et al., 2024) also leads to performance degradation at high sparsity ratios. Crucially, both methods fail to preserve client-specific knowledge and exacerbate conflicts during server-side aggregation under data heterogeneity. This results in an undesirable trade-off: reduced communication at the expense of significant performance loss. This leads us to our central research question:
@@
To address this question, we introduce FedSRD, a novel Sparsify-Reconstruct-Decompose framework for communication-efficient federated LLMs fine-tuning. Our framework is designed to maintain high model performance through three sequential operations.
@@ 2.3. Communication-Efficient Federated LLMs Finetuning
Communication overhead remains a primary bottleneck in FL systems (Yan et al., 2025, 2024; Sattler et al., 2019), a problem that is exacerbated when fine-tuning LLMs over the public internet. While PEFT methods provide a baseline level of efficiency, further optimizations are crucial. Structured methods impose constraints on the LoRA updates. For instance, FFA-LoRA (Sun et al., 2024) halves the communication cost by transmitting only the $B$ matrices. Fed-SB (Singhal et al., 2025) freezes the LoRA matrices entirely and instead inserts a small, trainable square matrix for aggregation. While these approaches reduce communication, they often do so at the cost of model performance by constraining the expressive capacity of LoRA. Other lines of work focus on generic compression. Quantization methods (Aji and Heafield, 2017; Konečnỳ et al., 2016) represent parameters with fewer bits, but they often fail to achieve high compression ratios and can lead to significant accuracy degradation, particularly in non-IID settings. Sparsification methods, which transmit only a subset of non-zero parameters, can achieve much higher compression ratios (Aji and Heafield, 2017; Sahu et al., 2021). Yet, traditional approaches are not directly applicable to LoRA. Magnitude-based pruning (Aji and Heafield, 2017) is suboptimal because it ignores the coupled relationship between the two LoRA matrices. Similarly, random sparsification (Yang et al., 2025; Yu et al., 2024) fails to preserve critical parameters, resulting in inferior performance at high sparsity ratios.
@@ 4.3. Alternative Decomposition
Previous works (Bai et al., 2024; Lee et al., 2025) often use Singular Value Decomposition (SVD) to decompose aggregated weights into new dense matrices $B^{t}$ and $A^{t}$. But transmitting them creates a download bottleneck. Naively sparsifying the full state matrices $B^{t}$ and $A^{t}$ would damage the learned model. Computing the update as a direct difference, $B^{t}-B^{t-1}$, is also mathematically flawed. This is because SVD is rotation-invariant and decomposition is not unique. Even a small global update can induce an arbitrary rotation of the low-rank basis between rounds, creating large, dense updates that do not reflect the true minimal updates. Thus, they are highly unsuitable for sparsification. These motivate the need for a novel decomposition capable of producing a stable update for sparse downloads.
@@ 5.3. Computation Time Analysis
We measured the wall-clock time required for a single FL round in our simulation environment for both the Llama3.2-3B and Qwen2-7B models, with results reported in Table 7. Among the baselines, FedAvg provides a baseline for time consumption. FedProx requires slightly more time due to its proximal term, whereas FFA-LoRA is faster as it only trains and aggregates half of the LoRA matrices. FlexLoRA consumes more time performing SVD for matrix decomposition, and FedICU is the most time-intensive due to its computationally inefficient disentangled aggregation. For our methods, while the full FedSRD requires more time for its $r$-rank SVD projection, our efficient variant, FedSRD-e, introduces negligible additional computation compared to FedAvg, while exhibiting performance second only to the full FedSRD. This result highlights that our FedSRD-e variant achieves a Pareto-optimal balance across model performance, communication efficiency, and time consumption, making it a highly practical and scalable solution for real-world federated LLMs fine-tuning.
@@ 6. Conclusion
In this paper, we address the critical challenge of communication overhead in federated LoRA fine-tuning for LLMs. We propose FedSRD, a novel Sparsify-Reconstruct-Decompose framework for communication-efficient federated LLMs fine-tuning. Its core components include a structure-aware, importance-based sparsification strategy on the client side; a conflict-mitigating reconstruction and aggregation step in the full-rank space on the server side; and a novel alternating decomposition method to produce sparse global updates for efficient download. We also introduce FedSRD-e, a computationally light variant for resource-constrained environments. Through extensive experiments on diverse benchmarks, we demonstrated that FedSRD reduces communication costs by over 90% while simultaneously improving model performance compared to state-of-the-art methods. Our work presents a practical and high-performing solution that significantly lowers the barrier for scalable, collaborative AI development on the decentralized Web.